two forces of magnitude 8N and 15N respectively act at a point.if the force is 17N,the angle between the forces has to be
1.60
2.45
3.90
4.30
1
Expert's answer
2012-10-26T11:53:01-0400
Let's write down the formula for the magnitude of the resultant:

R = √( a² + b² + 2*a*b*cos(α) ), (1)

where α is the angle between the forces. Therefore,

(1) ==> α = arccos (R² - a² - b²) / (2*a*b) =

= arccos (17² - 8² - 15²) / (2*8*15) =

= arccos 0 = π/2.

So, right answer is 3.90.